# Articulatory Encodec: Vocal Tract Kinematics as a Codec for Speech

### v1

@@ V Demonstration of Interpretability and Controllability of Articulatory Encodec
Unlike other high-dimensional neural coding, the proposed articulatory coding is highly interpretable and controllable. Interpretability is naturally granted by having a low-dimensional coding space, where each dimension corresponds to the physical articulator on the vocal tract. Furthermore, as the vocal tract articulation is the control space of speech production, the encoded articulatory features can be controlled with the same principle as how we produce speech. The control or modification of articulatory features then can be played back by the vocoder which functions as a physical simulation of the vocal tract articulation [3, 75, 12, 40, 16]. Here, we demonstrate these features with two cases.
@@ V-A Case 1: Place of Articulation – “lock” vs “rock”
We further demonstrate the simulated control by interpolating tongue articulations between the words [14]. The mixing factor, $\alpha$, is applied to weigh the sum of each of the tongue articulatory traces (TT, TB, and TD), i.e., $Art("lock")\cdot\alpha+Art("rock")\cdot(1-\alpha)$. The Fig. 5 bottom panel shows the interpolated traces with $\alpha\in\{1.0,0.8,0.6,0.4,0.2,0.0,-0.2\}$, gradually changing pronunciation from ”lock” to ”rock”. Then, ASR is applied to determine which word is perceived, drawing a perceptual boundary between the interpolation at $\alpha=0.2$, which is denoted as a separate line style in Fig. 5. This demonstration shows that the articulatory encodec provides interpretable control knobs, which can be manipulated to simulate speech sounds and observe the causal interaction between articulatory control and the generated speech.
@@ V-B Case 2: Manner of Articulation – Voice Onset Time
The next example is about the manner of articulation, especially focused on the phonological phenomena called voice onset time (VOT). While the spatial arrangement of articulators forms the filter that shapes the consonants, the manner of sound further categorizes the consonants into some perceptually distinguishable sounds. In this example, we demonstrate that the VOT can be directly implemented by the timing of the rise of loudness.
Here, a speech clip saying “bay” is encoded by the articulatory encoder (Fig. 6 left). We only show the traces of lips on the Y axis as these are the key articulatory factors of labial consonants. We can clearly see the lips opening around 200 ms. Then, we shift the loudness trace back and forth and generate speech from the manipulated features. By shifting the loudness trace 60 ms earlier than the moment of lips opening, we can nasalize the sound from “bay” to “may” (from plosive to nasal). On the other hand, by shifting it 60 ms backward, the sound is converted to “pay” (from lenis plosive to fortis plosive). We can also observe the induced VOT difference along this manipulation (Fig. 6 right), which is aligned with the known VOT patterns of nasal and plosive sounds.
These two proof-of-concept showcases suggest the promising utility of the articulatory encodec as a physical simulator of articulatory control, and as a speech analysis tool that provides a natural and intuitive phonological interpretation.
@@ V-C Potential Applications
We list some examples of potential applications here. First, the framework can be used as an analysis platform for investigating a phonological basis of speech without collecting high-cost articulatory data [16]. Second, the high-performance synthesizer can be used as a speech simulator, which can facilitate a control theoretic approach to speech processing or reinforcement learning of speaking agents [3, 76, 40, 77]. Third, the universal articulatory encoding can be utilized for a language learning tool or therapy, where looking through the vocal tract can aid the learning experience [78, 79, 80]. Lastly, a TTS system can be built upon the articulatory encodec that synthesizes articulatory control from text or higher-order structure of speech [81, 82, 83]. This can reveal a descriptive and interpretable relationship between text and articulation. To conclude, our framework has promising utilities in various domains in both speech science and engineering.
@@ VI Conclusion
We propose a novel encoding-decoding framework of speech, articulatory encodec, which mimics the biophysical interface of speech production. By projecting speech to articulatory features, the speech can be expressed in a low-dimensional space, where each channel corresponds to an articulator on the vocal tract. Such physical embodiment of speech coding enables a natural, intuitive interpretation, which is further facilitated by a visual aid that displays the articulation in a 2D animation. With large-scale training, our proposed articulatory encodec achieves a high-performance articulatory resynthesis that can generalize to unseen speakers and novel dialects, with a minimal loss of information compared to ground truth speech. To our knowledge, this is the first demonstration of universal articulatory synthesis that can scale up to an indefinite number of speakers. This universality is aided by a novel speaker identity encoding, which embeds highly discriminable speaker-specific speaker. Furthermore, the encoded speaker embedding is effectively disentangled from articulatory features and can be used for accent-preserving voice conversion.
In future work, we will scale the articulatory encodec to incorporate expressive speech and singing. Also, we will improve the robustness of the system under noisy environments. These efforts will maximize the promising utilities of the articulatory coding of speech.
@@ A.1. Layer-wise AAI Performance
We selected the 9-th layer after comparing the inversion performance of each layer measured by Pearson correlation coefficients (PCC) averaged across EMA channels on 5-fold cross-validation (Fig. 7). For each fold, 100 test utterances are randomly held out and 95 % confidence intervals are denoted. At the best layer (the 9th), the average correlation is $0.878\pm 0.012$. After the layer selection, all data from MNGU0 are used to fit the linear model.
@@
A.2. AAI Performance on Individual Articulator
@@
A.3. Evidence of Geometric Isometry Between Individual Articulatory Space
@@
A.4. Evidence for Excluding Coronal Axes of EMA

### v2

@@ V Demonstration of Interpretability and Controllability of Articulatory Encodec
Unlike other high-dimensional neural coding, the proposed articulatory coding is highly interpretable and controllable. Interpretability is naturally granted by having a low-dimensional coding space, where each dimension corresponds to the physical articulator on the vocal tract. Furthermore, as the vocal tract articulation is the control space of speech production, the encoded articulatory features can be controlled with the same principle as how we produce speech. The control or modification of articulatory features then can be played back by the vocoder, which functions as a physical simulation of the vocal tract articulation [3, 80, 12, 41, 16]. Here, we demonstrate these features with two example cases.
@@ V-A Case 1: Place of Articulation – “lock” vs “rock”
We demonstrate an articulatory control simulation by interpolating tongue articulations between the words [14]. The mixing factor, $\alpha$, is applied to weigh the sum of each of the tongue articulatory traces (TT, TB, and TD), i.e., $Art("lock")\cdot\alpha+Art("rock")\cdot(1-\alpha)$, where $Art(\cdot)$ means articulatory encoding. The Fig. 5 bottom panel shows the interpolated traces with $\alpha\in\{1.0,0.8,0.6,0.4,0.2,0.0,-0.2\}$, gradually changing pronunciation from ”lock” to ”rock”. Then, ASR is applied to synthesized speech to determine which word is perceived, drawing a perceptual boundary at $\alpha=0.2$, which is denoted using different line styles in Fig. 5. This demonstration shows that Articulatory Encodec provides interpretable control knobs, which can be manipulated to simulate speech sounds and observe the causal interaction between articulatory control and the generated speech.
@@ V-B Case 2: Manner of Articulation – Voice Onset Time
The next example is about the manner of articulation, especially focused on the phonological phenomena called voice onset time (VOT). While the spatial arrangement of articulators forms filters for consonants, the manner of sound further modifies sounds into perceptually distinguishable consonants. In this example, we demonstrate that the VOT can be directly implemented by the timing of the rise of loudness.
Here, a speech clip of the word “bay” is encoded by the articulatory encoder (Fig. 6 left). We only show the traces of lips on the Y axis as these are the key articulatory factors of labial consonants. We can clearly see the lips opening around 200 ms. Then, we shift the loudness trace back and forth and generate speech from the manipulated features. By shifting the loudness trace 60 ms earlier than the moment of lips opening, we can nasalize the sound from “bay” to “may” (from plosive to nasal). On the other hand, by shifting it 60 ms backward, the sound is converted to “pay” (from lenis plosive to fortis plosive). We can also observe the induced VOT difference along this manipulation (Fig. 6 right), which is aligned with the known VOT patterns of nasal and plosive sounds.
These two proof-of-concept showcases suggest the promising utility of Articulatory Encodec as a physical simulator of articulatory control, and as a speech analysis tool that provides a natural and intuitive phonological interpretation.
@@ V-C Potential Applications
Based on the results, we envision several potential applications of Articulatory Encodec. First, the framework can be used as an analysis platform for investigating a phonological basis of speech without collecting high-cost articulatory data [16]. Second, the high-performance synthesizer can be used as a speech simulator, which can facilitate a control theoretic approach to speech processing or reinforcement learning of speaking agents [3, 81, 41, 82]. Third, the universal articulatory encoding can be utilized for a language learning tool or therapy, where visualizing the vocal tract can enhance the learning experience [83, 84, 85]. Lastly, a TTS system can be built upon Articulatory Encodec that synthesizes articulatory control from text or higher-order structure of speech [86, 87, 88], revealing a descriptive and interpretable relationship between text and articulation. In conclusion, our framework holds promise for various applications in both speech science and engineering.
@@ VI Conclusion
We propose a novel encoding-decoding framework of speech, Articulatory Encodec, which mimics the biophysical interface of speech production. Through large-scale training, our proposed Articulatory Encodec framework achieves a high-performance articulatory inversion and synthesis that can generalize to unseen speakers, with a minimal loss of information compared to ground truth speech. To our knowledge, this is the first demonstration of universal articulatory synthesis that can scale up to an indefinite number of speakers. This universality is supported by a novel speaker identity encoding that embeds highly discriminable speaker-specific voice textures. Additionally, the encoded speaker embedding is effectively disentangled from articulatory features and can be used for accent-preserving voice conversion. Using the articulatory encodec, speech can be represented in a low-dimensional space, with each channel corresponding to an articulator in the vocal tract. This physical embodiment of speech allows for natural and intuitive control and interpretation.
In future work, we will scale the Articulatory Encodec framework to incorporate expressive speech and singing. Also, we will improve the robustness of the system under noisy environments. These efforts will maximize the promising utilities of the articulatory coding of speech.
@@ A.1. Layer-wise AAI Performance
We selected the 9-th layer after comparing the inversion performance of each layer measured by Pearson correlation coefficients (PCC) averaged across EMA channels on 5-fold cross-validation (Fig. 7). For each fold, 100 test utterances are randomly held out and 95 % confidence intervals are denoted. At the best layer (the 9th), the average correlation is $0.878\pm 0.012$. After the layer selection, all data from MNGU0 are used to fit the linear model. The parameters are estimated by ordinary least square (OLS) method. We find a minimal difference between with and without a regularization term (e.g., L2 penalty). Thus, we use OLS for a better simplicity.
@@
A.2. Measuring Correlation Between Articulations from Different Articulatory Space.
Since the EMA sensors are not spatially aligned across participants, each EMA channel from different participants is not directly comparable. As our model is using a particular participant’s (MNGU0’s) vocal tract as a template space, we find a mapping that spatially aligns MNGU0 sensors to different participants’ to calculate the correlation. Following [10], we pass speech audios from a target speaker from MOCHA-TIMIT or HPRC to our inversion model, and then a linear model is trained to predict the reference EMA of the target speaker. We used a linear regression with L1 penalty (Lasso) to impose sparsity in the weights, using $alpha=0.01$. Scikit-learn library is used for training the model (“sklearn.linear_model.Lasso”) [89]. This L1 penalty reveals channel-by-channel correspondence in the affine transformation (Fig. 9; more discussed in the next section). We split the data to 5 folds where 4 folds are used for training the linear model and 1 fold is used for evaluating the model. This is done individually for each speaker in MOCHA-TIMIT (7 speakers with 27 m of speech in average) and HPRC (7 speakers with 59 m of speech in average).
A.3. AAI Performance on Individual Articulator
@@
A.4. Evidence of Geometric Isometry Between Individual Articulatory Space
@@
A.5. Evidence for Excluding Coronal Axes of EMA